The wavelength of a greenlight from the traffic signal is Saturday at 5.20x10^-5cm. calculate the frequency
1
Expert's answer
2019-12-12T07:30:57-0500
"ν=\frac{c}{λ}=\frac{2.998\times 10^8ms^{-1}}{5.20\times10^{-7}m}= 5.77\times10^{14} s^{-1}"